OpenStudy (anonymous):

f(x)=5/x and g(x)=2x*2+5x what 2 numbers are not in the domain of the f o g

OpenStudy (amistre64):

fog would give us 5/g(x) when is g(x) = 0?

OpenStudy (anonymous):

i dont no im confused

OpenStudy (amistre64):

factoring g(x) is useful since we know that zero times anything is equal to zero 2x^2+5x ; factor out the common x to get x(2x+5) now, for what 2 values of x does: x(2x+5) = 0

OpenStudy (anonymous):

o ok thanks man

OpenStudy (amistre64):

can you tell me the values of x that would make x(2x+5) = 0 ?? keep in mind that a*b=0 when a=0 or b=0